# Exploiting Euclidean Distance Field Properties for Fast and Safe 3D planning with a modified Lazy Theta*

### v1

@@ 6 Conclusion and Future Work
A new heuristic search planner based on a modified Lazy Theta* planning algorithm that integrates EDF and exploits its properties has been presented. We proposed the FS-Planner that generates fast and safe paths, and eliminates the need for post-processing methods to smooth the paths. A new cost function has been proposed that adds a cost term related to the distance from obstacles. This provides a safe path planner.
The exploitation of the analytical properties of EDFs have enabled faster computation of the EDF costs along segments. We demonstrate the proposed approximation of $O(s_{i},s^{\prime}_{i+1})$ based on the analytical properties of the EDF (11).Thanks to the EDF properties can also reduce the number of visibility neighbours during path exploration from the obstacle-gradient information. Moreover, the way in which the selection of the visibility neighbors is performed also influences the safety of the path.
Heuristic search planners as Lazy Theta* algorithm take advantage of the triangle inequality to simplify the calculations to select the parent node. We have demonstrated that the EDF-based cost function used in the FS-Planner satisfies the triangle inequality and can be implemented in any heuristic algorithm such as A* or Theta*. Therefore, the FS-Planner improves with respect to CALT*+M1 because it does hold the triangle inequality, resulting in fewer calculations.
These claims have been validated with many tests and comparisons in 3D challenging scenarios and setups. One of the scenarios, S5, is real.
Future work will focus on the integration of new methods for online EDF computation based on neural radiation fields (NeRF). We will analyze how local path planners can take advantage of such online EDFs.

### v2

@@ 6 Conclusion and Future Work
A new heuristic search planner based on a modified Lazy Theta* planning algorithm that integrates EDF and exploits its properties has been presented. We proposed the FS-Planner, which generates fast and safe paths and eliminates the need for post-processing methods to smooth the paths. A new cost function has been introduced that incorporates a term related to the distance from obstacles, thus providing a safer path planner.
The exploitation of the analytical properties of EDFs have enabled faster computation of the EDF-based cost along segments. We demonstrate the proposed approximation of $O(s_{i},s^{\prime}_{i+1})$ based on the analytical properties of the EDF (11).Thanks to the EDF properties, the number of visibility neighbours can also be reduced during path exploration using the obstacle-gradient information. Moreover, the way in which visibility neighbours are selected directly influences the safety of the path.
Heuristic search planners as Lazy Theta* algorithm take advantage of the triangle inequality to simplify the calculations to select the parent node. We have demonstrated that the EDF-based cost function used in the FS-Planner satisfies the triangle inequality and can be implemented in any heuristic algorithm such as A* or Theta*. Therefore, the FS-Planner improves upon CALT*+M1 because it does hold the triangle inequality, resulting in fewer calculations.
These claims have been validated with extensive tests and comparisons in 3D challenging scenarios and setups. In addition to simulated indoor environments, we also evaluated the FS-Planner in three real-world outdoor scenarios, which provide complex and irregular obstacle configurations. These real-environment experiments complement the simulation study and further support the robustness and practical applicability of the FS-Planner.
Although the EDF-based cost term has been implemented here within a heuristic graph-search planner, the formulation itself is orthogonal to the underlying planning algorithm. Because the cost satisfies the triangle inequality and leverages analytical EDF properties, it can be incorporated into other planning paradigms, such as sampling-based methods, to potentially improve their safety and their ability to reason about obstacle proximity.
Overall, the FS-Planner deliberately trades strict theoretical guarantees of completeness and optimality for significant practical gains in computational efficiency. By reducing the number of explored visibility neighbours, the planner becomes sub-optimal and potentially incomplete; however, this relaxation leads to a structural reduction of the search space, which is particularly impactful in 3D environments with high branching factors. Although the fallback mechanism restores completeness by reverting to full neighbour exploration when required, it was never activated in any of the evaluated scenarios, including complex 3D and real-world environments. This observation should be interpreted as empirical evidence of robust behaviour rather than as a formal guarantee.
Furthermore, the benefits of the FS-Planner are not limited to computation time alone. The experimental results show that the reduction in computational cost, approximately $2\times$$3\times$, is accompanied by consistent improvements in geometric smoothness of the path, reflected in smaller heading changes, and by equal or higher safety levels in terms of obstacle clearance. Moreover, the FS-Planner offers an adjustable trade-off between efficiency and guarantees: configurations with fewer neighbours prioritise speed, while configurations with more neighbours approach optimal solutions. This flexibility makes the FS-Planner particularly suitable for local planning scenarios, where real-time performance, safety, and smoothnees are often more critical than strict optimality.
Future work will also include experiments with aerial robots, as well as the integration of new methods for online EDF computation based on neural radiation fields (NeRF). In addition, we will analyze how local path planners can take advantage of such online EDFs.